I am new to this subject and having difficulty in solving these problems. Please kindly help me.
1. Find three consequtive integers such that the sum of their squares is 77.3
2. solve equation w^3+5w^2-w=5
3. Factor completely z^4-81
Thank you for the help

Answer by stanbon(48546)

1. Find three consequtive integers such that the sum of their squares is 77.3
No solution. There are no three integers whose sum is 77.3
------------------
2. solve equation w^3+5w^2-w=5
w^3+5W^2-w-5 = 0
The coefficients sum to zero so 1 is a root and (x-1) is a factor.
Use synthetic division to find the quadratic factor:
1)....1....5....-1....-5
........1....6.....5...|..0
Remainder = 0
Quotient = x^2+6x+5 which factors as (x+5)(x+1)
--------------------
(w-1)(w+1)(w+5)=0
w=1 or w=-1 or w=-5
==============================
3. Factor completely z^4-81
= (z^4 - 3^4)
= (z^+9)(z^2-9)
= (z^2+9)(z+3)(z-3)
==================
Cheers,
Stan H.